The expression below is the factorization of what trinomial?
-1(x+6)(x+7)

Answer by ewatrrr(24785)

 
Hi,
-1(x+6)(x+7)= -(x^2 + 13x+42) = -x^2-13x-42
F First terms x^2
O Outside terms 7x
I Inside terms 6x
L Last terms 42